Question 1

Siphesihle borrows money from the bank at a discount rate of 16,5% per year. He must pay the bank R30 000 in eight months from now. The amount of money he receives from the bank now is

[1] R33 707,87.

[2] R26 700,00.

[3] R27 027,03.

[4] R33 463,26.

[5] R33 300,00.


Question 2

Phathu needs R10 500 in ten months' time to buy herself a new lens for her camera. Two months ago she deposited R9 000 in a savings account at a simple interest rate of 11,5% per year. How much money will Phathu still need to buy the lens ten months from now?

[1] R465,00

[2] R229,50

[3] R637,50

[4] R408,67

[5] None of the above


1
Expert's answer
2021-06-12T03:39:30-0400

Solution: for questions 1

Time, t = 8 months = 8/12 year = 2/3 year

Discount rate, r = 16.5% per year = 0.165

Let M be the borrowed money from the bank.

And D be the discounted amount.

Then, "D=M.r.t"

"\Rightarrow D=M(0.165)(2/3)
\\\Rightarrow D=0.11M ...(i)"

Now, "P=30000"

"P=M-D
\\ \Rightarrow 30000=M-0.11M" [Using (i)]

"\Rightarrow 0.89M=30000
\\\Rightarrow M=33,707.87"

Hence, option 1, i.e., R 33707.87 is correct.


For questions 2

Amount required after 10 months = R10,500

Two months ago amount deposited = R 9000

Interest Rate =11.5%

Interest rate per month (r)"= 11.5\%/12 = 0.958333 = 0.00958333"

Months of compounding for 9000 is 10+2 (n)= 12

Therefore, amount after 10 months "= 9000\times(1+r)^n"

"=9000\times(1+0.00958333)^{12}"

"=9000\times[1.12125928]"

"=10091.33"

Therefore, Amount required "=10500 - 10091.33 = 408.67"

The correct answer is  4.     R 408.67